OpenStudy (anonymous):

For events A and B, the following probabilities are known: Pr[A!B]=.8 andPr[A]=.5 What is the value of Pr[B]if A and B are disjoint (this means the events are partitions)?

OpenStudy (anonymous):

Any help would be great. An explanation would help

OpenStudy (kirbykirby):

is that Pr(A|B)=0.8?

OpenStudy (anonymous):

Pr[AUB] = .8 and Pr[A] = .5

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

trying to find Pr[b]

OpenStudy (anonymous):

and yes I believe

OpenStudy (kirbykirby):

Oh ok makes more sense now. \(P(A \cup B)=P(A)+P(B)-P(A \cap B)\) Since the events are disjoint, \(P(A\cap B)=0\) So \(P(B)=P(A\cup B)-P(A)\)

OpenStudy (anonymous):

ok so that is what I thought. it is equal to zero because it is disjoint. so really no resolution. correct?

OpenStudy (anonymous):

or that is the equation we have to work with? But yes that is what I am asking

OpenStudy (kirbykirby):

Um that is really just the formula for unions of events. And the intersection is 0 because A and B are disjoint. So you just solve for your missing variable, P(B).

OpenStudy (anonymous):

yes exactly which I can not figure out the method of finding the answer. having a hard time with this

OpenStudy (anonymous):

So P(B)=P(A∪B)−P(A) is our starting point so really are you saying .8-.5 would equal Pr(B)

OpenStudy (kirbykirby):

yes

OpenStudy (kirbykirby):

is that not the right answer?

OpenStudy (anonymous):

ahh you sure

OpenStudy (anonymous):

not sure. trying to find it out. that is what i had to work with

OpenStudy (kirbykirby):

Ok. well that should be the answer then :P

OpenStudy (anonymous):

wow so it really is 0

OpenStudy (anonymous):

thank you

OpenStudy (kirbykirby):

Yes. By definition of disjoint, \(P(A \cap B)=0\)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

make sense now

OpenStudy (kirbykirby):

:)